OpenStudy (anonymous):

Simplify: (5x - 10)2

OpenStudy (anonymous):

do you mean squared?

OpenStudy (anonymous):

\[25x ^{2}+100-100x\]

OpenStudy (mayliz):

25x^2+100-100x

OpenStudy (callisto):

(5x - 10)^2 = (5x)^2 - 2(5x)(10) +10^2 = 25x^2 - 100x +100

OpenStudy (callisto):

use the identity (a ± b)^2 = a^2 ± 2ab +b^2